The Bragg plates or crystalline material are separated by 1.2Armstrong. Calculate highest order Bragg reflection for 2.1ev
1
Expert's answer
2019-07-08T10:26:51-0400

Convert electron-volts to joules:


"E=2.1\text{ eV}\cdot1.602\cdot10^{-19}\frac{\text{J}}{\text{eV}}=3.36\cdot10^{-19}\text{ J}."

Hence calculate the wavelength:


"\lambda=\frac{hc}{E}=5.91\cdot10^{-7}\text{ m}."

Angstroms to meters:


"d=1.2\cdot10^{-10}\text{ m}."

Apply Bragg's law:


"2d\space\text{sin}\theta=n\lambda,"

express the order, which is a positive integer:


"n=\frac{2d\space\text{sin}\theta}{\lambda}=4.06\cdot10^{-4}\space\text{sin}\theta."

As we can see, with any value of the angle theta the order becomes close to 0 and by no means can become 1 for instance. Thus, we should choose a different wavelength for the experiment.